When 225 mL of H2O at 160C are mixed with 85 mL of water at 870C, what is the final temperature? (Assume that no heat is lost to the surroundings; dH2O = 1.00 g/mL. Cwater = 4.18 J/g. K )

Hint: the heat given by the hot water is received by the cold water.


1
Expert's answer
2022-01-07T16:27:01-0500

q=mc∆t

 For water q1 = q2

 m1c∆t1 = m2c∆t2

If the density of water is 1.00 g/mL then m1= 225 g and m2 = 85 g 

 225 × (tf - 16) = 85 × (87 - tf)

 tf = 35.4677

Final temperature= 35.4677 degrees celsius.